A harmonic oscillator has a mass of 4 kg and a spring with a force constant 100

N/m. Find the period, frequency and angular frequency.


1
Expert's answer
2022-02-28T10:41:14-0500

The period

"T=2\pi\sqrt{\frac{m}{k}}=2\pi\sqrt{\frac{4}{100}}=0.4\pi=1.26\:\rm s"

The frequency

"f=1/T=1/1.26=0.8\:\rm Hz"

The angular frequency

"\omega=2\pi f=5\:\rm rad/s"